If a polynomial f(x) is such that f(0)=2, f(1)=3, f(x+2)=2f(x)-f(x+1), then find f(5)

Grade:7

1 Answers

Arun
4 years ago
Dear student
 
put x= 0 in equation
f(2) = 2 f(0) - f(1) = 4-3 = 1
 
Now put x = 1
f(3) = 2 f(1) – f(2) = 6-1 = 5
 
now put x = 2
f(4) = 2f(2) – f(3) = 2 – 5 = – 3
 
Now put x = 3
f(5) = 2f(3) – f(4) = 10 + 3 = 13
 
Hope it helps